A normally distributed data has a mean of 200 and a standard deviation of 20.What is the lowest value that belongs to the upper 10% of the distribution?

1
Expert's answer
2022-04-28T16:03:07-0400
"P(X>x)=1-P(Z\le\dfrac{x-\mu}{\sigma})"

"=1-P(Z\le\dfrac{x-200}{20})=0.10"

"\dfrac{x-200}{20}\approx1.281552"

"x=200+20(1.281552)"

"x=225.631"